# Exact analytical algorithm for solvent accessible surface area and derivatives in implicit solvent molecular simulations on GPUs

### v1

@@ 3.1 Molecular SASA estimation
In the test set to validate SASA estimation, we selected eight proteins from the set of the previously examined proteins for ab initio protein folding 2, in which the set of proteins have diverse topologies: Trp-cage (20 residues), Fip35 (33 residues), NTL9 (39 residues), BBL (47 residues), NuG2variant (56 residues), Cspa (69 residues), Lambda-repressor (80 residues) and Top7 (92 residues). An ensemble of structures for every protein were extracted from the protein folding trajectories in that work with an even interval, obtaining a set of conformations with diverse atomic and molecular SASA values. The corresponding data were estimated for every conformation using ICOSA, LCPO and pwSASA and dSASA. The numerical method ICOSA can provide more accurate SASA estimation for molecules than LCPO and pwSASA, so we mainly compared the molecular SASA values from dSASA with ICOSA. The results are shown in Figure 6 and every point represents one conformation in the trajectory. In all the systems, dSASA can generate well-correlated SASA values with ICOSA as Pearson correlation coefficients of the linear regression are $R^{2}=0.98,0.99$ and the slopes are also close to 1. We note that this close agreement in fact demonstrates the relative accuracy of ICOSA as compared with our exact geometric method.
@@
The molecular SASA values from LCPO and pwSASA were calculated using the same conformations and compared with ICOSA method. As shown in Figure 7, the discrepancy of LCPO molecular SASA values with ICOSA values varies for different conformations in the systems. Overall, the LCPO tends to underestimate the values with most of the values are above the perfect agreement line. The correlation coefficients are around 0.9 but it is lower than 0.9 for NTL9 and Cspa while the slopes are less than 0.9 for these two systems as well. The results became worse with pwSASA shown in Figure 8, where the similar phenomenon arose: the discrepancy of pwSASA molecular SASA values with ICOSA values varies for conformations in the systems; furthermore, all the correlation coefficients are less than 0.9 and the coefficients can be less than 0.8 for NTL9, BBL, NuG2variant and Cspa.
@@
Both LCPO and pwSASA use pairwise overlaps which can be relatively faster, but the inexact calculation will lead to inaccuracy especially for some types of atoms, for example both methods occasionally produce negative SASA values for individual atoms while the atomic values from dSASA are all nonnegative. LCPO includes the higher order correction terms to account for the overlap between two neighbors of the central atom. Then several atom types are defined by the environment, such as the atomic number, the number of bonded neighbors and the state of hybridization. Every atom type has predefined parameters through training. Because of less data for certain atom types, the parameter values for these atoms may generate unexpected values which explains the discrepancy for the proteins above. As noticed in 28, the method works better for more exposed atoms and tends to overestimate the surface area of some buried atoms, such as the carboxyl carbons of the amino acids Asp, Asn, Gln and Glu. pwSASA only considers the overlap between neighboring pairs and was trained based on a number of protein systems to match the values of LCPO. Even though introducing more atom types, it suffers similar problems as LCPO, likely due to insufficient training data and accounting for 3-body and higher terms only in an average way. In fact, the pwSASA values are not the original SASA values because it introduces an empirical adjustment to the SASA values to compensate for a systematic divergence from the numerical ICOSA calculations, which may adversely affect the overall accuracy as well.
@@ 3.2 Speed Comparison in GB/SA MD
The original version of dSASA was implemented on CPUs, and the transformation to GPU version speeds up the computation by taking advantage of the parallel computing. The program written in CUDA was integrated to the Amber20 version. In sander, pmemd, or pmemd.cuda, setting the $gbsa$ flag to 4 in GB suite will activate GB/SA simulations with dSASA. We first compared the speed of GB/SA MD on GPU with CPU and next with other methods using 10 proteins with various sizes ranging from 10 residues to 190 residues. 4 of them were selected from above and the PDB code of the rest are given: 10 (PDB code, 5AWL), 20, 33, 52 (2P6J), 69, 92, 110 (1BYW), 130 (1E6K), 162 (2B75) and 190 (1BK7).
We examined the method on the CPU Intel Platinum 8268 2.90GHz and on the GPU Nvidia RTX 2080 Ti. The estimated wallclock time for one step SASA calculation on CPU is from 12 ms (10 residues) to 90 ms (190 residues), while the time trend on GPU is from 12 ms to 30 ms. The wallclock time here and below is averaged over several simulations. The speed on CPU is comparable to the speed on GPU when the size of protein is small, and the speed on GPU becomes relatively faster as the size of the protein increases. The estimated wallclock time in hours required to complete 1 ns GB/SA simulation in Amber is given in the left of Figure 9. The wallclock time required to finish 1 ns simulation on the CPU increases from 0.87 hour (10 residues) to 33.80 hours (190 residues), while the corresponding trend of timing on the GPU is from 0.6 hour to to 1.87 hours. When the size of protein is small, the improvement of GPU over CPU version is small with multiplicity 1.44 times for 10 residues, while the multiplicity grows as the size of molecule is increasing with 18.07 times for 190 residues.
@@
We next compared the speed with other methods in Amber using the same set of systems above. The wallclock time in hours needed to complete 1 ns simulation is provided in the right of Fig. 9. The speed of pwSASA is stable and much faster than other two methods as it is designed to have a fast and simple estimation with more approximations, so we can confirm that the accuracy of the method is at cost of speed as LCPO and dSASA can reproduce more accurate SASA values by the testing above. As can be seen in Fig. 9, the speed of dSASA is more stable as the size of the system grows, while LCPO has a steeper increasing trend when the size of the molecule becomes larger. When the size of the system is small (10 residues), dSASA is slower than LCPO. While the crossing point is around the size of 69 residues and after that, dSASA starts to gain advantage over LCPO for 69, 92 residues and the multiplicity becomes greater for larger systems. This phenomenon fits the computation complexity for these two methods: the computation complexity of LCPO is $O(N^{2})$ while the complexity of dSASA is $O(N\log N)$ (the complexity of tetrahedrization algorithm), so the system with a larger size will have relatively better performance. Another reason is that the GB/SA simulation with LCPO is CPU/GPU hybrid so the data transfer further hinders the performance. On the other hand, pwSASA is $\sim$80 times faster than dSASA for small proteins, reducing to $\sim$50 times for larger proteins. While pwSASA is a pairwise method and its computation complexity is also $O(N^{2})$, its computation is efficiently embedded in the overall energy computation which is of the same order.
@@ 3.3 Stability Analysis of Proteins
dSASA can estimate the molecular SASA accurately in proteins with diverse shapes based on the testing above. With the speedup on GPUs, it is now possible to examine its performance in GB/SA simulations and evaluate its effect on the stability of protein structures. Then we simulated two proteins with dSASA and compared the results with experiments, GB-only and other SASA methods. The simulation results are still dominated by the force field and the solvent model, so such comparisons should be used by caution as the inaccuracies in the solute force field may affect the simulations and the solvent models may become less reliable away from 300 K. The trends in the simulations can provide useful information albeit with these limitations. The thermal stability profiles of Trp-cage and the fraction of folded on homeodomain are calculated from REMD at various temperatures. The temperature ladder for Trp-cage contains 8 replicas (247.7-387.3 K, see Methods for full list). The ladder for homeodomain has 10 replicas (288.7-407.8 K).
@@
As investigated in 20, the surface tension 5 cal/($\cdot$ Å${}^{2}$) is working properly and is the default value in GB/SA in Amber, and larger value 7 cal/($\cdot$ Å${}^{2}$) is expected to give more structures within native range and the value is 7.2 cal/($\cdot$ Å${}^{2}$) as proposed in 5. For Trp-cage system, we tested the surface tension 5 and 7 to check the performance of the method. For homeodomain, we only used the value 7, which can produce more near-native states in the trajectories.
@@
The thermal stability profiles of Trp-cage with surface tension 5 and 7 cal/($\cdot$ Å${}^{2}$) were first computed and the fractions of the near native conformations are given in Figure 10 for all methods. The left shows the results with the surface tension 5 cal/($\cdot$ Å${}^{2}$), while the right shows the results with the surface tension 7 cal/($\cdot$ Å${}^{2}$). At the surface tension 5, all three methods incorporating the nonpolar term can achieve better agreement with experiments by creating closer thermal stability profiles at various temperatures. At 300 K, dSASA can generate $\sim$75fraction of folded states while other methods with nonpolar term and experimental results are $\sim$80and the GB-only can produce $\sim$50%. When the temperature is greater than 310 K, the fraction of folded from dSASA is still slightly smaller than other methods but dSASA can produce closer results to the experiments. The predicted melting temperature $T_{m}$ is 316.8 K and the experimental result is around 317 K, which is better than the GB-only (predicted $T_{m}$ is around 300 K) and other methods (predicted $T_{m}$ is greater than 320 K for LCPO and pwSASA). Moreover, it is expected that a larger surface tension can generate more native-like conformations through which we can prove that the nonpolar term does not destroy the system. As shown in the right of Figure 10, all three methods can generate more native-like structures with surface tension 7 cal/($\cdot$ Å${}^{2}$).
@@
Figure 10: Thermal stability profiles for Trp-cage with surface tension 5 (left) and 7 cal/($\cdot$ Å${}^{2}$)(right) in GB and GB/SA simulations respectively, including experimental data.
@@
To show the changing process of RMSD, we provide the detailed trajectories at several temperatures for Trp-cage starting from extended and native states with surface tension 5 cal/($\cdot$ Å${}^{2}$) in Supporting Information.
@@
For homeodomain variant, we first simulated REMD with surface tension 5 cal/($\cdot$ Å${}^{2}$) using pwSASA for 2 $\mu s$, but we noticed that the fraction of folded was far from the experimental results at higher temperatures. The experimental measured $T_{m}$ is greater than 372 K, so we expect a larger surface tension can achieve better agreement with experiments. We then simulated with surface tension 7 cal/($\cdot$ Å${}^{2}$) for all GB/SA methods. All three methods with nonpolar term can improve the fraction of folded at various temperatures compared to GB-only. The fractions of native-like starting from native and unfolded states at selected temperatures are shown in Figure 11. Since the experimental $T_{m}$ is greater than 372 K, we expect that the system will not collapse at high temperatures. As two better methods, LCPO and dSASA can produce more native-like conformations at higher temperatures as shown in Figure 11. Moreover, dSASA generated more native-like conformations than LCPO at various temperatures, proving that our method can have a better effect on stabilizing the system. The detailed trajectories of homeodomain at selected temperatures starting from extended and native states with surface tension 7 cal/($\cdot$ Å${}^{2}$) are given in Supporting Information. Furthermore, at the temperature 300 K, the difference between the native-like fractions of the simulations starting from native and extended is around 20% for dSASA and LCPO after 500 ns simulations, while that difference is around 40% for GB and pwSASA even after 1000 ns simulations, indicating that the inclusion of an accurate SASA term might be able to result in converging simulations at a faster rate. The improved performance from dSASA still has variance from the experimental results, for example the fraction of native-like at 377.7 K is around 35% while it is 50% in experiment. Such discrepancy may come from limitations of the SASA-based nonpolar term, or from the force field and the polar GB term, which are under active development.
We carried out another set of simulations for LCPO and dSASA, and the trajectories of fraction of folded are given in Supporting Information Figure 18. The simulation with dSASA can produce relatively less native-like conformations at lower temperatures in the trajectories starting from the native state while the fractions of folded starting from extended do not increase much at higher temperatures. On the other hand, in LCPO, the trajectories starting from native state produce less native-like conformations at lower temperatures than the trajectories starting from extended state. The results from both methods were not as expected because the MD simulations converge more slowly as the topological complexity of the system increases. Longer time simulations are clearly needed in such situation.
@@ 3.4 Performance on RNA systems
Another advantage of dSASA over pwSASA is that it can be applied to other types of systems such as nucleic acids or small-molecule ligands, because only the radii of the atoms are considered without defining diverse atom types. Here we test this aspect using two RNA systems, 2KOC and 1ESH, and compare the molecular SASA values with ICOSA and other methods to see if the excellent performance on proteins is transferable to other biopolymers.
[…]
@@ 4 Conclusions
In this work, we present dSASA, an analytical SASA evaluation method for molecules and its implementation on GPUs. In this approach, the weighted Delaunay tetrahedrization is first computed with the atomic coordinates and radii. Next the atomic and molecular SASA values are estimated using inclusion-exclusion method based on exterior simplices from tetrahedrization, resulting in more accurate estimation than LCPO and pwSASA in reproducing numerical ICOSA SASA values. The pre-trained parameters for atoms in LCPO and pwSASA can speed up the calculation, but the data sometimes will provide biased values for certain types of atoms so the computed molecular SASA values had smaller correlation coefficients with the values from ICOSA method, such as in the protein systems Cspa and Nug2variant and RNAs. The estimation from dSASA can reproduce ICOSA values with correlation coefficients greater than 98%. Moreover, the current GPU version speeds up the calculation substantially compared with the CPU version, making it applicable for longer time GB/SA MD simulations especially for larger systems. The gain of speed arises from the complexity of the algorithm itself which will scales well for larger systems and from the full implementation on GPU devices removing the data transfer with CPUs. dSASA had a stable trend of speed as the size of molecule increases but it was still relatively slower than pwSASA. However, dSASA started to outperform LCPO when the system contains around 70 residues, and the performance for larger molecules would be more stable than LCPO because the computation in LCPO is pairwise and its implementation is CPU/GPU hybrid in Amber.
@@
In the GB/SA simulations, two proteins (Trp-cage and homeodomain) were simulated and compared to other methods and experimental results. dSASA achieved comparable performance with LCPO and pwSASA on the small protein, and the performance became better for the larger system. The simulated melting curve with nonpolar term for Trp-cage was more consistent with the experimental measures compared with that without nonpolar term. In homeodomain variant, the melting temperature is greater than 372 K so we expect the system can maintain some amount of native-like conformations at higher temperatures. The simulations generated more extended conformations with other methods at high temperatures while dSASA can produce more native-like conformations in the trajectories which is closer to the experimental data.
@@
As the program has been rigorously examined, we anticipate to extend its applications to other functions in Amber software, such as post-processing of the trajectories. Given the accuracy of dSASA, the calculation of atomic and molecular SASA values can be benchmark data set for the training of the parameters in pwSASA approach for RNAs in the future, which will combine the advantage of its speed on GPUs and the accurate calculation from dSASA. The surface tension was assumed to be identical for all atoms in the implementation. As the SASA value is exact for every atom in dSASA, we can attempt to assign different surface tension values to atoms according to the atomic environment which is believed to be able to achieve more accurate solvation free energies. Furthermore, the calculation of SASA is based on the diagram of Laguerre intersection cells, so it can be easily extended to the computation of molecular volumes along with the corresponding atomic derivatives. Inclusion of the volume term will provide a more complete description for the nonpolar solvation term. With the volume derivatives, we can further examine the impact of the term on the stability of molecules in the MD simulations. Moreover, the current algorithm being used for weighted Delaunay Tetrahedrization, gReg3D, is designed to work on a set of random points and the size of the workspace depends on the distribution of points in the workspace. As it consumes nearly 70% of wallclock time of our surface area calculation, improvement on this algorithm will further speed up the simulations. The spatial information of the atoms and the molecular properties in Amber were not fully utilized, then the influence of including such information among atoms could be further explored. The program of dSASA written in CUDA will be freely available from the authors.

### v2

@@ 3.1 Molecular SASA estimation
In the test set to validate SASA estimation, we selected eight proteins from the set of the previously examined proteins for ab initio protein folding 2, in which the set of proteins have diverse topologies: Trp-cage (20 residues), Fip35 (33 residues), NTL9 (39 residues), BBL (47 residues), NuG2variant (56 residues), CSPA (69 residues), Lambda-repressor (80 residues) and Top7 (92 residues). An ensemble of structures for every protein were extracted from the protein folding trajectories in that work with an even interval, obtaining a set of conformations with diverse atomic and molecular SASA values. The corresponding data were estimated for every conformation using ICOSA, LCPO and pwSASA and dSASA. The numerical method ICOSA can provide more accurate SASA estimation for molecules than LCPO and pwSASA, so we mainly compared the molecular SASA values from dSASA with ICOSA. The results are shown in Figure 6 and every point represents one conformation in the trajectory. In all the systems, dSASA can generate well-correlated SASA values with ICOSA as Pearson correlation coefficients of the linear regression are $R^{2}=0.98,0.99$ and the slopes are also close to 1. We note that this close agreement in fact demonstrates the relative accuracy of ICOSA as compared with our exact geometric method. However, ICOSA lacks the derivatives needed for MD simulation.
@@
The molecular SASA values from LCPO and pwSASA were calculated using the same conformations and compared with ICOSA method. As shown in Figure 7, the discrepancy of LCPO molecular SASA values with ICOSA values varies for different conformations in the systems. Overall, the LCPO tends to underestimate the values with most of the values falling above the perfect agreement line. The correlation coefficients are around 0.9, but the slope and correlation are are lower than 0.9 for NTL9 and CSPA. The results became worse with pwSASA shown in Figure 8, where the similar phenomenon arose: the discrepancy of pwSASA molecular SASA values with ICOSA values varies for conformations in the systems; furthermore, all the correlation coefficients are less than 0.9 and the coefficients can be less than 0.8 for NTL9, BBL, NuG2variant and CSPA. This reflects on the goal of pwSASA being speed rather than accuracy.
@@
Both LCPO and pwSASA use pairwise overlaps which can be relatively faster, but the inexact calculation will lead to inaccuracy especially for some types of atoms. For example, both methods occasionally produce unphysical negative SASA values for individual atoms, while the atomic values from dSASA are all nonnegative. LCPO includes the higher order correction terms to account for the overlap between two neighbors of the central atom. Then several atom types are defined by the environment, such as the atomic number, the number of bonded neighbors and the state of hybridization. Every atom type has predefined parameters through training. Because of less data for certain atom types, the parameter values for these atoms may generate unexpected values which explains the discrepancy for the proteins above. As noticed in 28, the method works better for more exposed atoms and tends to overestimate the surface area of some buried atoms, such as the carboxyl carbons of the amino acids Asp, Asn, Gln and Glu. pwSASA only considers the overlap between neighboring pairs and compensated by training highly specific parameters applicable only to the atomic environments present in standard amino acids. Even though pwSASA introduced many additional empirical parameters, it suffers similar problems as LCPO, likely due to weak transferability of the pair parameters, and accounting for 3-body and higher terms only in an average way. In fact, the pwSASA values are not the original SASA values because it introduces an empirical adjustment to the SASA values to compensate for a systematic divergence from the numerical ICOSA calculations, which may adversely affect the overall accuracy as well.
@@ 3.2 Speed Comparison in GB/SA MD
The original version of dSASA was implemented on CPUs, and the transformation to a GPU version speeds up the computation by taking advantage of the parallel computing. The program written in CUDA was integrated to the Amber20 version. In sander, pmemd, or pmemd.cuda, setting the $gbsa$ flag to 4 in GB suite will activate GB/SA simulations with dSASA. We first compared the speed of GB/SA MD on GPU with CPU and next with other methods using 10 proteins with various sizes ranging from 10 residues to 190 residues. 4 of them were selected from above and the PDB code of the rest are given: 10 (PDB code, 5AWL), 20, 33, 52 (2P6J), 69, 92, 110 (1BYW), 130 (1E6K), 162 (2B75) and 190 (1BK7).
We benchmarked the method on the CPU Intel Platinum 8268 2.90GHz and on the GPU Nvidia RTX 2080 Ti. The estimated wallclock time for one step SASA calculation on CPU is from 12 ms (10 residues) to 90 ms (190 residues), while the time trend on GPU is from 12 ms to 30 ms. The wallclock time here and below is averaged over several simulations. The speed on CPU is comparable to the speed on GPU when the size of protein is small, and the speed on GPU becomes relatively faster as the size of the protein increases. The estimated wallclock time in hours required to complete 1 ns GB/SA simulation in Amber is given in the left of Figure 9. The wallclock time required to finish 1 ns simulation on the CPU increases from 0.87 hour (10 residues) to 33.80 hours (190 residues), while the corresponding trend of timing on the GPU is from 0.6 hour to to 1.87 hours. When the size of protein is small, the improvement of GPU over CPU version is small with speedup 1.44 times for 10 residues, while the speedup grows as the size of molecule is increasing to 18.07 times for 190 residues. For the system with 190 residues, we tested the parallel version with multiple CPUs. The GPU version is still much faster than the parallelized CPU version. The speedup of the GPU version over parallel version with 4, 8, 16 CPUs is 7.6, 9.1, 4.8, respectively.
@@
We next compared the speed with other methods in Amber using the same set of systems above. The wallclock time in hours needed to complete 1 ns simulation is provided in the right of Fig. 9. The speed of pwSASA is stable and much faster than other two methods as it is designed to have a fast and simple estimation with more approximations, so we can confirm that the accuracy of the method is at cost of speed as LCPO and dSASA can reproduce more accurate SASA values by the testing above. As can be seen in Fig. 9, the speed of dSASA is more stable as the size of the system grows, while LCPO has a steeper increasing trend when the size of the molecule becomes larger. When the size of the system is small (10 residues), dSASA is slower than LCPO. While the crossing point is around the size of 69 residues and after that, dSASA starts to gain advantage over LCPO for 69, 92 residues and the speedup becomes greater for larger systems. This phenomenon fits the computation complexity for these two methods: the computation complexity of LCPO is $O(N^{2})$ while the complexity of dSASA is $O(N\log N)$ (the complexity of tetrahedrization algorithm), so the system with a larger size will have relatively better performance. Another reason is that the GB/SA simulation with LCPO is CPU/GPU hybrid so the data transfer further hinders the performance. On the other hand, pwSASA is $\sim$80 times faster than dSASA for small proteins, reducing to $\sim$50 times for larger proteins. While pwSASA is a pairwise method and its computation complexity is also $O(N^{2})$, however its computation is efficiently embedded in the overall energy computation which is of the same order; thus pwSASA adds little to the overall computational cost.
@@ 3.3 Stability Analysis of Proteins
dSASA can estimate the molecular SASA accurately in proteins with diverse shapes based on the testing above. With the speedup on GPUs, it is now possible to examine its performance in GB/SA simulations and evaluate its effect on the stability of protein structures. However, comparison of the simulation results to experiments should be done with the understanding that deviations can arise from inaccuracy in any of the multiple components of the model (such as the biopolymer force field or GB model, which typically contribute more than the SASA term). Additionally the implicit solvation models may become less reliable away from 300 K. Given these caveats, we simulated two proteins with dSASA and compared the results with experiments, GB-only and other SASA methods. The trends in the simulations can provide useful information albeit with these limitations. The thermal stability profiles of Trp-cage and the fraction of folded on homeodomain are calculated from REMD at various temperatures. The temperature ladder for Trp-cage contains 8 replicas (247.7-387.3 K, see Methods for full list). The ladder for homeodomain has 10 replicas (288.7-407.8 K).
@@
The default surface tension value in Amber GB/SA module is 5 cal/($\cdot$ Å2) 4 and the recommended value is 7 cal/($\cdot$ Å2) 5 for pwSASA 20. For Trp-cage, we tested surface tension values of 5 and 7 to examine the performance of the method. For homeodomain, we only used the value 7, which can produce more near-native states in the trajectories.
@@
The thermal stability profiles of Trp-cage with surface tension 5 and 7 cal/($\cdot$ Å2) were first computed and the fractions of the near native conformations are given in Figure 10 for all methods. The left shows the results with the surface tension 5 cal/($\cdot$ Å2), while the right shows the results with the surface tension 7 cal/($\cdot$ Å2). With surface tension of 5, all three methods incorporating the nonpolar term achieve better agreement with experiments by creating more stable thermal stability profiles at various temperatures. At 300 K, simulations using dSASA generate $\sim$75fraction of folded states while other methods with nonpolar term and experimental results are $\sim$80and the GB-only leads to $\sim$50%. When the temperature is greater than 310 K, the fraction of folded from dSASA is still slightly smaller than other methods but dSASA can produce closer results to the experiments. The predicted melting temperature $T_{m}$ is 316.8 K and the experimental result is around 317 K, which is better than the GB-only (predicted $T_{m}$ is around 300 K) and other methods (predicted $T_{m}$ is greater than 320 K for LCPO and pwSASA). Moreover, it is expected that a larger surface tension can generate more native-like conformations through which we can demonstrate that the nonpolar term does not disrupt the system. As shown in the right of Figure 10, all three methods can generate more native-like structures with surface tension 7 cal/($\cdot$ Å2).
@@
Figure 10: Thermal stability profiles for Trp-cage with surface tension 5 (left) and 7 cal/($\cdot$ Å2)(right) in GB and GB/SA simulations respectively, including experimental data.
@@
To show the extent of dynamics and variation in sampled RMSD, we provide the detailed trajectories at several temperatures for Trp-cage starting from extended and native states with surface tension 5 cal/($\cdot$ Å2) in Supporting Information.
@@
For the homeodomain variant, we first carried out REMD with surface tension 5 cal/($\cdot$ Å2) using pwSASA for 2 $\mu s$, but we noticed that the fraction of folded was far from the experimental results at higher temperatures. The experimental measured $T_{m}$ is greater than 372 K, so we expect a larger surface tension can achieve better agreement with experiments. We then simulated with surface tension 7 cal/($\cdot$ Å2) for all GB/SA methods. All three methods with nonpolar term can improve the fraction of folded at various temperatures compared to GB-only. The fractions of native-like starting from native and unfolded states at selected temperatures are shown in Figure 11. Since the experimental $T_{m}$ is greater than 372 K, we expect that the system will not collapse at high temperatures. The two more accurate SASA methods, LCPO and dSASA, produce more native-like conformations at higher temperatures as shown in Figure 11. Moreover, dSASA generated more native-like conformations than LCPO at various temperatures, demonstrating that our method can have a better effect on stabilizing the system. The detailed trajectories of homeodomain at selected temperatures starting from extended and native states with surface tension 7 cal/($\cdot$ Å2) are given in Supporting Information. Furthermore, at the temperature 300 K, the difference between the native-like fractions of the simulations starting from native and extended is around 20% for dSASA and LCPO after 500 ns simulations, while that difference is around 40% for GB and pwSASA even after 1000 ns simulations, indicating that the inclusion of an accurate SASA term might lead to improved sampling and faster convergence (or conversely, that the SASA approximations may lead to increased kinetic traps). The improved performance from dSASA still has variance from the experimental results, for example the fraction of native-like at 377.7 K is around 35% while it is 50% in experiment. Such discrepancy may come from limitations of the SASA-based nonpolar term, or from the force field and the polar GB term; comparison of simulation results to experiment typically does not reveal which component is responsible, and further model system development is needed.
@@ 3.4 Performance on RNA systems
Another advantage of dSASA over pwSASA is that it can be applied to other types of systems such as nucleic acids or small-molecule ligands, because only the radii of the atoms are considered without defining diverse atom types. Here we test this aspect using two RNA systems, 2KOC and 1ESH, and compare the molecular SASA values with ICOSA and other methods to see if the excellent performance on proteins is transferable to other biopolymers. Simulation results for RNA are not presented in details, since accurate force fields and implicit solvent models for RNA are still a matter of active research 45, 46.
To validate the SASA estimation, we generated alternate conformations for these RNA systems with the GB-only method for 1 $\mu s$ and extracted 104 conformations for each system respectively. The comparison of the molecular SASA values between all methods is shown in Figure 12. dSASA can reproduce SASA values from ICOSA with more than 98% accuracy for both systems while LCPO and pwSASA produced divergent results for two systems and they had overall underestimation of the SASA values. The SASA estimates from pwSASA have large discrepancy from ICOSA method, likely because it was trained exclusively on protein data so that the parameters do not accurately model the specific atom types and local environments present in RNA. Overall, the results confirm the expected outcome that dSASA accuracy is transferable across diverse molecular systems.
@@ 4 Conclusions
In this work, we present dSASA, an analytical SASA evaluation method for molecules and its implementation on GPUs. In this approach, the weighted Delaunay tetrahedrization is first computed with the atomic coordinates and radii. Next the atomic and molecular SASA values are estimated using inclusion-exclusion method based on exterior simplices from tetrahedrization, resulting in more accurate estimation than LCPO and pwSASA in reproducing numerical ICOSA SASA values. The pre-trained parameters for atoms in LCPO and pwSASA can speed up the calculation, but these parameters can have low transferability. This is confirmed by comparison of computed molecular SASA values, which for these approximate methods had smaller correlation coefficients with the values from ICOSA method, such as in the protein systems CSPA and NuG2variant and RNAs. In contrast, estimation from dSASA can reproduce ICOSA values with correlation coefficients greater than 98%. Moreover, the current GPU version speeds up the calculation substantially compared with the CPU version, making it applicable for longer time GB/SA MD simulations especially for larger systems. The gain of speed arises from the complexity of the algorithm itself which will scales well for larger systems and from the full implementation on GPU devices removing the data transfer with CPUs. dSASA has a stable trend of speed as the size of molecule increases but it is still relatively slower than the highly approximate pwSASA. However, dSASA started to outperform LCPO when the system contains around 70 residues, and the performance for larger molecules would be more stable than LCPO because the computation in LCPO is pairwise and its implementation is CPU/GPU hybrid in Amber.
@@
In the GB/SA simulations, two proteins (Trp-cage and homeodomain) were simulated and compared to other methods and experimental results. dSASA achieved comparable performance with LCPO and pwSASA on the small protein, and the performance became better for the larger system. The simulated melting curve with nonpolar term for Trp-cage was more consistent with the experimental measures compared with that without nonpolar term. In homeodomain variant, the melting temperature is greater than 372 K so we expect the system can maintain some amount of native-like conformations at higher temperatures. The simulations generated more extended conformations with other methods at high temperatures while dSASA can produce more native-like conformations in the trajectories which is closer to the experimental data. However, these comparisons to experiment must be considered in the context of the other force field components as well.
@@
As the program has been rigorously examined, we anticipate to extend its applications to other functions in Amber software, such as post-processing of the trajectories in MM-PBSA or MM-GBSA. The calculation in dSASA is geometry based, obtaining accurate results with the provided atomic coordinates and radii. Therefore it can be used for diverse types of systems without restrictions, such as the simulations of larger proteins, protein-ligand complexes, and protein-nucleic acids complexes. 1 nanosecond GB/SA simulation for 200-residue proteins with dSASA takes around two hours. Even though it is slower than pwSASA, the accuracy of the method will allow us to explore the impact of the nonpolar term on the simulations in the future. Given the accuracy of dSASA, the calculation of atomic and molecular SASA values can be benchmark data set for the training of the parameters in pwSASA approach for RNAs in the future, which will combine the advantage of its speed on GPUs and the accurate calculation from dSASA. The SASA is certainly an approximation to nonpolar solvation, but many studies have shown that including it improves agreement with experiment for things like protein stability or binding affinities12. In addition, dSASA, a more accurate SASA with derivatives implemented on GPUs, will also enable fast polar solvation methods for MD. Furthermore, the calculation of SASA is based on the diagram of Laguerre intersection cells, so it can be easily extended to the computation of molecular volumes along with the corresponding atomic derivatives. Inclusion of the volume term will provide a more complete description for the nonpolar solvation term. With the volume derivatives, we can further examine the impact of the term on the stability of molecules in the MD simulations. Moreover, the current algorithm being used for weighted Delaunay Tetrahedrization, gReg3D, is designed to work on a set of random points and the size of the workspace depends on the distribution of points in the workspace. As it consumes nearly 70% of wallclock time of our surface area calculation, improvement on this algorithm will further speed up the simulations. The program of dSASA written in CUDA will be freely available from the authors.
5 Associated Content
Supporting Information available.
Examination of energy conservation after including dSASA term in MD. Detailed trajectories with dSASA, GB, LCPO and pwSASA for Trp-cage and Homeodomain at selected temperatures. Fractions of folded for the second test set of Homeodomain.